OpenStudy (anonymous):

Could someone help me simplify this radical \[\sqrt{225x^13yz^6}\]

OpenStudy (anonymous):

suppose to be x to the 13th power

OpenStudy (zzr0ck3r):

\[=\sqrt{25*9*3*x^{13}*y*z^6}=\sqrt{25}*\sqrt{9}*\sqrt{3}*\sqrt{x^{13}}*\sqrt{y}*\sqrt{z^6}=\] \[5*3*\sqrt{3}*\sqrt{x^{13}}*\sqrt{y}*\sqrt{z^6}=15x^6z^3*\sqrt{3xy}\]

OpenStudy (anonymous):

I think you have an extra factor of three under the radical.\[\sqrt{225x^{13} yz^6}=15x^6z^3\sqrt{xy}\]Otherwise, my solution follows zzr0ck3r.